A 50.0 kg person is walking horizontally with a constant acceleration of 0.25 m/s2 inside an elevator. The elevator is also accelerating downward at a rate of 1.0 m/s2. Find the force of friction between his shoes and the floor of the elevator. What minimum force of friction between his shoes and the floor would make this scenario possible?
1
Expert's answer
2019-04-26T11:27:05-0400

The normal force is:


"N=mg-ma_1=m(g-a_1)=50(9.8-1)=440\ N"

The force of friction between his shoes and the floor of the elevator:


"F_f=ma_2=(50)(0.25)=12.5\ N"

The minimum coefficient of friction between his shoes and the floor:


"\mu =\frac{F_f}{N}=\frac{12.5}{440}=0.028"